Please help Algebra 2 assignment ?
Show work
There’s a red line straight through the 60 degree radian/3
1) draw a right triangle at the given angle. Label sides x and y
2) what special right triangle is this similar to? What is the ratio of the sides of this special right triangle?
3) Use the ratio to find the side lengths of the triangle(values of x and y)
4) translate the values to all four quadrants
1
Expert's answer
2020-01-08T13:55:50-0500

Answer:


1.






2.It is 90°-60°-30° triangle, "1: \sqrt{3} : 2" is the ratio of lengths of sides.


3.Since Z = z than "X =z\/2" and "Y=\sqrt{3}*z/2"


4.In first quadrant: "X =z\/2" and "Y=\sqrt{3}*z/2"


In second quadrant: "X = - z\/2" and "Y=\sqrt{3}*z/2"


In third quadrant: "X =-z\/2" and "Y=-\sqrt{3}*z/2"


In fourth quadrant:"X =z\/2" and "Y=-\sqrt{3}*z/2"